OVHCloud Network Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
GSW <> BHS
Incident Report for Network & Infrastructure
Resolved
We have 4 links down between our GlobalSwitch (Paris) and Beauharnois (Canada) sites. We've warned our provider.

Update(s):

Date: 2017-07-13 06:23:50 UTC
Outage Stop: July 13, 2017 00:30 GMT

Date: 2017-07-12 20:35:58 UTC
Reparation have started. The new ETA given by our provider is 23:30 UTC.

Date: 2017-07-12 19:27:01 UTC
The repair team has some trouble reaching the site. New ETA is 19:30 UTC.

Date: 2017-07-12 18:24:57 UTC
We got an ETR from our provider: 18:30 UTC. We will check as soon as the link are up again.

Date: 2017-07-12 17:02:16 UTC
The following comment come from our provider.

\"Event Field Services have performed fiber testing procedures and have isolated the suspected fiber cut 43 kilometers from the testing site. Field Services are en route to the location to further isolate the damaged fiber and formulate the repair plan.\"

They did not provide us with an ETR yet. Thank you for your patience.

Date: 2017-07-12 15:31:35 UTC
Incident should be resolved by 16:30 UTC.

Date: 2017-07-12 14:24:16 UTC
Our provider is informing us that there is a fiber cut in Hammonton, NJ. We are waiting for more information.

Date: 2017-07-12 14:23:21 UTC
A fiber cut in Hammonton, NJ is the cause of these links down. We are waiting for an ETR from our provider.
Posted Jul 12, 2017 - 14:04 UTC